OpenStudy (anonymous):

Find the values of X.Y.Z 2x+y+z=5 x-y+2z=4 x+y+z=4

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

@Hero help

OpenStudy (lgbasallote):

add equation 2 and equation 3

OpenStudy (anonymous):

how?

OpenStudy (lgbasallote):

x - y + 2z = 4 x + y + z = 4 what's x + x?

OpenStudy (anonymous):

x^2?

OpenStudy (lgbasallote):

nope that is x times x

OpenStudy (anonymous):

2x?

OpenStudy (lgbasallote):

better

OpenStudy (lgbasallote):

now what's -y + y

OpenStudy (anonymous):

0

OpenStudy (lgbasallote):

good..what's 2z + z?

OpenStudy (anonymous):

3z

OpenStudy (lgbasallote):

and 4+4?

OpenStudy (anonymous):

8

OpenStudy (lgbasallote):

therefore x -y + 2z = 4 x + y + z = 4 ------------ 2x + 3z = 8 <---let this be equation 4 (we set it aside for now)

OpenStudy (lgbasallote):

now add equation 1 and 2

OpenStudy (anonymous):

okay i did that and i got 3x+3x=9 do i combine that with equation 4? @igbasllote

OpenStudy (lgbasallote):

dont you mean 3x + 3z = 9?

OpenStudy (lgbasallote):

coz im reading 3x + 3x = 9..my eyes could be blurred though

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

well then very good you combine it with equation 4 3x + 3z = 9 2x + 3x = 8 the question is...do you add or subtract?

OpenStudy (anonymous):

add?? :/

OpenStudy (lgbasallote):

let's try... 3x + 3z = 9 2x + 3x = 8 ----------- 5x + 6z = 17 <---you still have two variables remember the point of this thing is to lessen the number of variables...

OpenStudy (lgbasallote):

so since adding didnt work you subtract

OpenStudy (anonymous):

ahh okay

OpenStudy (lgbasallote):

so subtract 3x + 3x = 9 2x + 3x = 8 what do you get as difference?

OpenStudy (anonymous):

1x or X and 0z and 1

OpenStudy (anonymous):

(@lgbasallote lol you have been typing 2x + 3x =8, think it's supposed to 2x + 3z = 8)

OpenStudy (anonymous):

@nphuongsun93 right! lol

OpenStudy (lgbasallote):

lol im making the same mistake =)))))

OpenStudy (lgbasallote):

anyway @sharee2012 you're right x = 1

OpenStudy (anonymous):

ahh sleep calls lol

OpenStudy (anonymous):

thank you :D very much for the help

OpenStudy (lgbasallote):

now substitute x = 1 into equation 4 or 5 whichever you prefer to get z...then substitute x and z into equation 1,2 or 3

OpenStudy (lgbasallote):

thank you for being a participative learner <tips hat>

OpenStudy (anonymous):

im a little confused

OpenStudy (lgbasallote):

confused where?

OpenStudy (anonymous):

the pluggin in to solve the last part -_-

OpenStudy (lgbasallote):

have you found out what z is yet?

OpenStudy (anonymous):

is z 0?

OpenStudy (anonymous):

@igbasallote?

OpenStudy (anonymous):

Not an i, it's an L. :) Z is not zero. 3x+3z = 9 You've already found x = 1, so substitute the value of x: 3(1)+3z=9

OpenStudy (lgbasallote):

lol sorry i was not able to come...you tagged the wrong account it's LGBASALLOTE. IGBASALLOTE is a different user lol

OpenStudy (lgbasallote):

anyway thanks for coming on my behalf @Wired

OpenStudy (anonymous):

oh okay sorry im so tired lol and okay i get it @Wired yes thank u very much :D

OpenStudy (anonymous):

No problemo